### v1

# S1-NexusAgent: a Self-Evolving Agent Framework for Multidisciplinary Scientific Research

Abstract
Modern scientific research relies on large-scale data, complex workflows, and specialized tools, which existing LLMs and tool-based agents struggle to handle due to limitations in long-horizon planning, robust goal maintenance, and continual learning from execution. To address these issues, in this work, we propose S1-NexusAgent, a self-evolving agent framework designed for multidisciplinary scientific research. S1-NexusAgent adopts a hierarchical Plan-and-CodeAct execution paradigm, decoupling global scientific planning from subtask-level tool execution through a dual-loop architecture, thereby enabling stable modeling of complex research workflows. The system natively supports the Model Context Protocol (MCP), integrates up to thousands of cross-disciplinary scientific tools, and achieves efficient orchestration of heterogeneous research tools via intention-aware dynamic tool retrieval and hot-plug mechanisms. To address long-context and large-scale data challenges in scientific settings, S1-NexusAgent introduces object-reference-based sparse context management, which enables sub-task context isolation and intermediate result compression. Building on this, a Critic Agent automatically evaluates complete execution trajectories and distills high-quality research paths into reusable Scientific Skills, forming a closed loop for continuous self-evolution, which is valuable for sustainable and long-horizon scientific research. Experiments on authoritative scientific benchmarks involving long-horizon planning and complex specialized tool orchestration, including biomini-eval (biology), ChemBench (chemistry), and MatSciBench (material science), demonstrate that S1-NexusAgent achieves state-of-the-art performance, validating its effectiveness and generalization capability in complex scientific tasks.
1 Related Work
1.1 Domain-Specific Scientific Agents
A large body of recent work builds domain-specialized scientific agents that encode substantial task priors for a particular field, such as molecular design and optimization, materials discovery/characterization, and biomedical discovery workflows. Representative systems include chemistry and materials agents like ChemCrow [Bran et al., 2024], CACTUS [McNaughton et al., 2024], IR-Agent [Noh et al., 2025], HoneyComb [Zhang et al., 2024], MatPilot [Ni et al., 2024], MAPPS [Zhou et al., 2025], and others, as well as life-science agents spanning drug discovery, genomics/bioinformatics, and protein/CRISPR applications (e.g., DrugAgent [Inoue et al., 2024]/DrugPilot [Li et al., 2025a], Biomni [Huang et al., 2025], BioDiscoveryAgent [Roohani et al., 2024], CellVoyager [Alber et al., 2025], GeneGPT [Jin et al., 2024], CRISPR-GPT [Huang et al., 2024], etc.). These agents often achieve strong performance by aligning planning and tool use with a narrowly defined workflow or domain toolchain.
While domain-specific agents can be highly effective within a fixed scope, S1-NexusAgent targets cross-domain scientific tasks with a unified architecture, emphasizing long-horizon stability, flexible tool orchestration over a large heterogeneous tool ecosystem, and system-level mechanisms (e.g., dual-loop execution, context structuring, and self-evolution) that are not tied to a single domain’s workflow.
1.2 Complex General Scientific Agent Systems
Beyond domain-specific scientific agents, a small line of work targets multidisciplinary scientific agent systems that aim to support scientific problem solving across multiple disciplines within a unified framework. Unlike agents designed for a single field or workflow, these systems emphasize domain-agnostic architectural mechanisms, relying on dynamic tool use rather than fixed pipelines to adapt to different scientific contexts. Representative examples include InternAgent [Zhang et al., 2025a], which coordinates heterogeneous agents and tools to handle cross-domain scientific tasks, and SciMaster [Chai et al., 2025], which explores general-purpose scientific intelligence through scalable, tool-augmented agentic workflows evaluated across diverse scientific domains, such as materials science [Zhang et al., 2025c], physics [Miao et al., 2025], machine learning [Zhu et al., 2026], etc.
While InternAgent and SciMaster highlight the feasibility of multidisciplinary scientific agents, they largely rely on monolithic or loosely structured control loops. In contrast, S1-NexusAgent explicitly introduces an inner–outer dual-loop architecture that separates long-horizon, task-level planning from subtask-level executable reasoning. This separation enables S1-NexusAgent to maintain global goal stability while supporting localized trial-and-error, systematic replanning, and trajectory-level evaluation, providing a clearer structural basis for long-term robustness and self-evolution through reusable Scientific Skills.
1.3 Code-enhanced and programmatic scientific agents
Another line of related work emphasizes code as a first-class action and planning medium for scientific agents. Rather than relying solely on natural-language reasoning, these systems generate executable artifacts—such as programs, scripts, or structured workflows—to drive scientific computation, data analysis, and experimental validation, improving precision, reproducibility, and controllability. Representative systems include AlphaEvolve [Novikov et al., 2025], which performs iterative program synthesis and evaluation for scientific optimization; workflow-oriented agents such as AIGS [Liu et al., 2024] and ORGANA [Darvish et al., 2025], which synthesize executable pipelines for scientific processes; and tool-augmented analysis agents such as Chemist-X [Chen et al., ] and Biomni [Huang et al., 2025], which iteratively generate and execute code to validate hypotheses and refine results. In these systems, planning is often encoded implicitly through program structure or explicitly via intermediate program representations, with execution serving as the primary grounding mechanism.
S1-NexusAgent aligns with the code-centric philosophy of these approaches but differs in key aspects. It adopts CodeAct [Wang et al., 2024] as a unified execution interface, explicitly coupling reasoning and code execution within an iterative inner loop rather than treating code generation as a one-off artifact. Moreover, code execution is embedded within an explicit inner–outer dual-loop architecture, where programmatic actions support local hypothesis validation while global planning and strategy revision occur at the task level. Finally, S1-NexusAgent incorporates trajectory evaluation and Scientific Skill distillation, enabling successful code-driven executions to be accumulated and reused—extending code-enhanced agents from execute-to-solve toward execute-to-learn-and-reuse in long-horizon scientific workflows.
2 Architecture
3 Agentic Training
4 Quantitative evaluation
5 Case studies of S1-NexusAgent
6 Discussion
To rigorously evaluate the individual contributions of our methodology, we conduct systematic ablation studies along two complementary dimensions, with all experiments evaluated on the Biomni Eval-1 benchmark: (1) the effectiveness of the proposed Inner–Outer Dual-Loop architecture, and (2) the impact of staged training strategies, comparing supervised fine-tuning (SFT) and reinforcement learning (RL), on agentic execution quality.
For the Dual-Loop architecture ablation, all architectural variants are evaluated on Biomni-Eval-1 using Qwen3-8B-SciCodeAct—a model specifically fine-tuned for scientific programming—as a fixed backbone, ensuring that performance differences are solely attributable to architectural design.
For the training strategy ablation, we likewise conduct evaluation on Biomni-Eval-1, fixing the backbone to the base Qwen3-8B model and comparing Only SFT against SFT + RL under identical conditions, thereby isolating the effect of reinforcement learning on agentic execution.
6.1 Ablation on the Dual-Loop Architecture
Scientific research tasks impose dual requirements on an autonomous agent: maintaining long-horizon strategic coherence while retaining flexibility for local, tool-driven exploration. The proposed Dual-Loop architecture explicitly separates these concerns by decoupling task-level planning (Outer Loop) from execution-level exploration and error recovery (Inner Loop).
To isolate the contribution of each component, we compare the full S1-NexusAgent against two degenerate variants: Only Outer-loop, where execution is limited to a single-pass CodeAct call without iterative feedback, and Only Inner-loop, where global task decomposition and replanning are disabled.
As summarized in Table 1, we observe the following findings:
  • Critical Role of Local Exploration. The Only Inner-loop configuration (38.99% average) outperforms the Only Outer-loop variant (37.23% average), indicating that iterative CodeAct-driven exploration and error recovery constitute a non-trivial contributor to performance in complex biological tasks.
  • Synergistic Integration of Planning and Execution. The complete Dual-loop system achieves the best overall performance (42.42% average), suggesting that the Outer Loop functions as a corrective mechanism that mitigates suboptimal local execution and stabilizes long-horizon reasoning, particularly in tasks such as Rare Disease Diagnosis.
Table 1: Ablation study of the Dual-Loop Architecture on Biomni-Eval-1. Results compare isolated loop configurations against the integrated framework. Best results are highlighted in bold.
Framework Overall Patient CRISPR Lab-Bench Lab-Bench Rare GWAS GWAS GWAS GWAS Screen
Design Avg. Gene Delivery SeqQA DbQA Disease Catalog OT $P$ VP Design
Only Outer-loop 37.23 8.0 40.0 46.0 26.0 10.0 38.0 74.0 70.0 16.3 44.0
Only Inner-loop 38.99 14.0 40.0 44.0 24.0 13.3 40.0 80.0 74.0 18.6 42.0
Dual-loop 42.42 18.0 60.0 42.0 30.0 23.3 50.0 72.0 72.0 20.9 36.0
6.2 Ablation on RL-based Staged Training
We further analyze how staged learning enhances the execution quality of the CodeAct module. In this context, staged training refers to an SFT initialization that enforces protocol adherence, followed by RL-based policy optimization that refines decision-making under uncertainty. Using the same Qwen3-8B backbone, we compare Only SFT against SFT + RL.
Key observations from Table 2 include:
  • Substantial Gains from RL-based Optimization. Incorporating RL yields a marked improvement in overall success rates (42.42% vs. 38.40% for SFT alone), with especially pronounced gains on tasks characterized by complex decision trees, such as GWAS Catalog (50.0% vs. 24.0%).
  • Beyond Format Compliance. While SFT primarily enforces structured tool usage, RL enables the agent to acquire more robust research strategies by leveraging on-policy execution feedback, allowing it to better cope with tool-specific idiosyncrasies and distribution shifts inherent in multidisciplinary scientific workflows.
Table 2: Ablation study of RL-based Staged Training on Biomni-Eval-1. All configurations utilize the Qwen3-8B-SciCodeAct backbone. Best results are highlighted in bold.
Training Overall Patient CRISPR Lab-Bench Lab-Bench Rare GWAS GWAS GWAS GWAS Screen
Stage Avg. Gene Delivery (SeqQA) (DbQA) Disease (Catalog) (OT) ($P$) (VP) Design
Only SFT 38.40 18.0 60.0 34.0 28.0 20.0 24.0 80.0 72.0 14.0 34.0
SFT + RL 42.42 18.0 60.0 42.0 30.0 23.3 50.0 72.0 72.0 20.9 36.0
Notably, the SFT + RL configuration corresponds to the full Dual-loop NexusSciAgent, indicating that architectural synergy and training strategy jointly contribute to the observed performance gains.
7 Conclusion
In this work, we present S1-NexusAgent, a self-evolving agent framework designed for long-horizon, tool-intensive, and multidisciplinary scientific research. By adopting a hierarchical Plan-and-CodeAct paradigm, S1-NexusAgent explicitly decouples global scientific planning from subtask-level execution through an inner–outer dual-loop architecture, enabling stable goal maintenance while supporting localized exploration and trial-and-error. The framework further introduces an extensible scientific tool ecosystem with native MCP support, intent-aware dynamic tool retrieval, and object-reference-based sparse context management, collectively addressing scalability challenges arising from large tool collections, long contexts, and data-intensive workflows.
Beyond architectural design, S1-NexusAgent incorporates a trajectory-evaluation-based self-evolution mechanism, where a Critic Agent systematically evaluates complete execution trajectories and distills high-quality scientific behaviors into reusable Scientific Skills. This design enables the system to accumulate experience across tasks and domains, forming a closed loop of execution, evaluation, and capability growth. We further demonstrate that agentic post-training of the inner-loop CodeAct module—via supervised fine-tuning and reinforcement learning—significantly improves execution robustness and scientific tool interaction, even when using relatively small open-weight models.
Extensive experiments on Biomni-Eval, ChemBench, and MatSciBench show that S1-NexusAgent achieves state-of-the-art performance across life sciences, chemistry, and materials science, validating both its effectiveness and generalization capability in complex scientific tasks. Case studies further illustrate how the dual-loop architecture and code-centric execution paradigm enable interpretable, adaptive, and resilient scientific workflows in realistic research scenarios.
Overall, S1-NexusAgent represents a step toward scalable, adaptive, and continuously improving scientific agents, bridging the gap between static tool-augmented assistants and autonomous systems capable of long-term scientific reasoning and learning. We believe this framework provides a strong foundation for future research on agentic learning, cross-domain scientific skill transfer, and AI systems that more closely resemble the iterative and reflective processes of human scientists.
Team Members
  • First Authors:
    Ziliang Pang (CASIA)
    *Shuo Ren (CASIA)
  • Core Contributors:
    Zhang Zhou (CASIA)
    Xiaomian Kang (CASIA)
    Qianlong Du (CASIA)
    Can Xie (CASIA)
    Ziyi Zhang (CASIA)
    Zhenjiang Ren (CASIA)
  • Contributors (listed in alphabetical order by last name):
    Yingying Chen (CASIA)
    Guangchuan Guo (Taichu)
    Longteng Guo (CASIA)
    Wenjing Li (CASIA)
    Zhiyuan Li (CASIA)
    Dan Liu (CASIA)
    Xiaotao Wang (CASIA)
    Yuying Wang (Taichu)
    Chunpeng Yang (Taichu)
    Ge Yang (CASIA)
    Haisheng Yu (Taichu)
    Zhidong Zhang (Taichu)
    Yang Zhao (CASIA)
    Bingke Zhu (CASIA)
  • Supervisors (listed in alphabetical order by last name):
    Jian Cheng (CASIA)
    Rufeng Chu (Taichu)
    Linjing Li (CASIA)
    Jing Liu (CASIA)
    Jinqiao Wang (CASIA)
    Ge Yang (CASIA)
    Zhaoxiang Zhang (CASIA)
    Dajun Zeng(CASIA)
    Chengqing Zong (CASIA)
  • Project Lead:
    *Jiajun Zhang (CASIA)
    (*corresponding authors)
CASIA: Institute of Automation, Chinese Academy of Sciences.
Taichu: Zhongke Zidong Taichu (Beijing) Technology Co., Ltd.

### v2

@@
 Experiments on authoritative scientific benchmarks involving long-horizon planning and complex specialized tool orchestration, including biomini-eval (biology), ChemBench (chemistry), and MatSciBench (material science), demonstrate that S1-NexusAgent achieves state-of-the-art performance, validating its effectiveness and generalization capability in complex scientific tasks.
+The code of S1-NexusAgent is available at https://github.com/CASIA-LM/S1-NexusAgent.
 1 Related Work
 1.1 Domain-Specific Scientific Agents
-A large body of recent work builds domain-specialized scientific agents that encode substantial task priors for a particular field, such as molecular design and optimization, materials discovery/characterization, and biomedical discovery workflows.
+A large body of recent work builds domain-specialized scientific agents that encode substantial task priors for a particular field, such as molecular design and optimization, materials discovery/characterization, and biomedical discovery workflows [Ren et al., 2026].
 Representative systems include chemistry and materials agents like ChemCrow [Bran et al., 2024], CACTUS [McNaughton et al., 2024], IR-Agent [Noh et al., 2025], HoneyComb [Zhang et al., 2024], MatPilot [Ni et al., 2024], MAPPS [Zhou et al., 2025], and others, as well as life-science agents spanning drug discovery, genomics/bioinformatics, and protein/CRISPR applications (e.g., DrugAgent [Inoue et al., 2024]/DrugPilot [Li et al., 2025a], Biomni [Huang et al., 2025], BioDiscoveryAgent [Roohani et al., 2024], CellVoyager [Alber et al., 2025], GeneGPT [Jin et al., 2024], CRISPR-GPT [Huang et al., 2024], etc.).
@@
 Rather than relying solely on natural-language reasoning, these systems generate executable artifacts—such as programs, scripts, or structured workflows—to drive scientific computation, data analysis, and experimental validation, improving precision, reproducibility, and controllability.
-Representative systems include AlphaEvolve [Novikov et al., 2025], which performs iterative program synthesis and evaluation for scientific optimization; workflow-oriented agents such as AIGS [Liu et al., 2024] and ORGANA [Darvish et al., 2025], which synthesize executable pipelines for scientific processes; and tool-augmented analysis agents such as Chemist-X [Chen et al., ] and Biomni [Huang et al., 2025], which iteratively generate and execute code to validate hypotheses and refine results.
+Representative systems include AlphaEvolve [Novikov et al., 2025], which performs iterative program synthesis and evaluation for scientific optimization; workflow-oriented agents such as AIGS [Liu et al., 2024] and ORGANA [Darvish et al., 2025], which synthesize executable pipelines for scientific processes; and tool-augmented analysis agents such as Chemist-X [Chen et al., 2023] and Biomni [Huang et al., 2025], which iteratively generate and execute code to validate hypotheses and refine results.
 In these systems, planning is often encoded implicitly through program structure or explicitly via intermediate program representations, with execution serving as the primary grounding mechanism.
@@
 6 Discussion
-To rigorously evaluate the individual contributions of our methodology, we conduct systematic ablation studies along two complementary dimensions, with all experiments evaluated on the Biomni Eval-1 benchmark: (1) the effectiveness of the proposed Inner–Outer Dual-Loop architecture, and (2) the impact of staged training strategies, comparing supervised fine-tuning (SFT) and reinforcement learning (RL), on agentic execution quality.
-For the Dual-Loop architecture ablation, all architectural variants are evaluated on Biomni-Eval-1 using Qwen3-8B-SciCodeAct—a model specifically fine-tuned for scientific programming—as a fixed backbone, ensuring that performance differences are solely attributable to architectural design.
-For the training strategy ablation, we likewise conduct evaluation on Biomni-Eval-1, fixing the backbone to the base Qwen3-8B model and comparing Only SFT against SFT + RL under identical conditions, thereby isolating the effect of reinforcement learning on agentic execution.
+To rigorously evaluate the individual contributions of our methodology, we conduct systematic ablation studies along two complementary dimensions, with all experiments evaluated on the Biomni-Eval1 benchmark: (1) the effectiveness of the proposed Inner–Outer Dual-Loop architecture, and (2) the impact of staged training strategies, comparing supervised fine-tuning (SFT) and reinforcement learning (RL), on agentic execution quality.
+For the Dual-Loop architecture ablation, all architectural variants are evaluated on Biomni-Eval1 using Qwen3-8B-Sci—a model specifically fine-tuned for scientific programming—as a fixed backbone, ensuring that performance differences are solely attributable to architectural design.
+For the training strategy ablation, we likewise conduct evaluation on Biomni-Eval1, fixing the backbone to the base Qwen3-8B model and comparing Only SFT against SFT + RL under identical conditions, thereby isolating the effect of reinforcement learning on agentic execution.
 6.1 Ablation on the Dual-Loop Architecture
@@
 The complete Dual-loop system achieves the best overall performance (42.42% average), suggesting that the Outer Loop functions as a corrective mechanism that mitigates suboptimal local execution and stabilizes long-horizon reasoning, particularly in tasks such as Rare Disease Diagnosis.
-Table 1: Ablation study of the Dual-Loop Architecture on Biomni-Eval-1. Results compare isolated loop configurations against the integrated framework. Best results are highlighted in bold.
+Table 1: Ablation study of the Dual-Loop Architecture on Biomni-Eval1. Results compare isolated loop configurations against the integrated framework. Best results are highlighted in bold.
 Framework Overall Patient CRISPR Lab-Bench Lab-Bench Rare GWAS GWAS GWAS GWAS Screen
@@
 In this context, staged training refers to an SFT initialization that enforces protocol adherence, followed by RL-based policy optimization that refines decision-making under uncertainty.
-Using the same Qwen3-8B backbone, we compare Only SFT against SFT + RL.
+Using the same Qwen3-8B backbone, we compare the baseline with Only SFT and SFT + RL.
 Key observations from Table 2 include:
@@
 While SFT primarily enforces structured tool usage, RL enables the agent to acquire more robust research strategies by leveraging on-policy execution feedback, allowing it to better cope with tool-specific idiosyncrasies and distribution shifts inherent in multidisciplinary scientific workflows.
-Table 2: Ablation study of RL-based Staged Training on Biomni-Eval-1. All configurations utilize the Qwen3-8B-SciCodeAct backbone. Best results are highlighted in bold.
+Table 2: Ablation study of RL-based Staged Training on Biomni-Eval1. Best results are highlighted in bold. The baseline takes Qwen-8B as the backbone without post-training.
 Training Overall Patient CRISPR Lab-Bench Lab-Bench Rare GWAS GWAS GWAS GWAS Screen
 Stage Avg. Gene Delivery (SeqQA) (DbQA) Disease (Catalog) (OT) ($P$) (VP) Design
+Baseline 33.63 10.0 60.0 32.0 24.0 3.3 16.0 76.0 74.0 7.0 34.0
 Only SFT 38.40 18.0 60.0 34.0 28.0 20.0 24.0 80.0 72.0 14.0 34.0
@@
 Notably, the SFT + RL configuration corresponds to the full Dual-loop NexusSciAgent, indicating that architectural synergy and training strategy jointly contribute to the observed performance gains.
+These results indicate that, in multi-step, multi-tool scientific task settings, the execution layer (the inner-loop in our case) is not merely a passive reasoning endpoint, but a highly learnable and adaptable core component.
+Through agentic training applied solely to the execution layer, the model achieves substantial improvements in both execution stability and task efficiency, thereby significantly enhancing the reliability of reasoning and task execution for scientific agents in complex scenarios.
+This observation further motivates an important research direction: given that complex scientific tasks typically involve long-horizon decision trajectories with extended execution times, it may be a key avenue for advancing the overall capabilities of scientific agents to explore joint optimization mechanisms between the execution layer and the strategic layer within hierarchical agent architectures, to enable cross-level cooperative learning and long-term decision consistency.
 7 Conclusion
@@
 We further demonstrate that agentic post-training of the inner-loop CodeAct module—via supervised fine-tuning and reinforcement learning—significantly improves execution robustness and scientific tool interaction, even when using relatively small open-weight models.
-Extensive experiments on Biomni-Eval, ChemBench, and MatSciBench show that S1-NexusAgent achieves state-of-the-art performance across life sciences, chemistry, and materials science, validating both its effectiveness and generalization capability in complex scientific tasks.
+Extensive experiments on Biomni-Eval1, ChemBench, and MatSciBench show that S1-NexusAgent achieves state-of-the-art performance across life sciences, chemistry, and materials science, validating both its effectiveness and generalization capability in complex scientific tasks.
 Case studies further illustrate how the dual-loop architecture and code-centric execution paradigm enable interpretable, adaptive, and resilient scientific workflows in realistic research scenarios.